there is a promotion where the customer selects a ball from a ballot box if a red one is selected not for the account. It is known that in the ballot box there are 99 black balls and a red one. There are ten people trained in the row. What is the probability? that two people win the promotion?
Answer #1

Probability of selecting red one, p = 1/100 = 0.01

q = 1 - p

Number of people, n = 10

This is a case of binomial distribution. P(X) = nCx px qn-x

P(2 people win the promotion) = P(X = 2)

= 10C2 x 0.012 x 0.998

= 0.00415